# Constrained Decoding for Robotics Foundation Models constrained-robot-fms.github.io

### v1

@@ Abstract
Recent advances in the development of robotic foundation models have led to promising end-to-end and general-purpose capabilities in robotic systems. These models are pretrained on vast datasets of robot trajectories to process multimodal inputs and directly output a sequence of action that the system then executes in the real world. Although this approach is attractive from the perspective of improved generalization across diverse tasks, these models are still data-driven and, therefore, lack explicit notions of behavioral correctness and safety constraints. We address these limitations by introducing a constrained decoding framework for robotics foundation models that enforces logical constraints on action trajectories in dynamical systems. Our method ensures that generated actions provably satisfy signal temporal logic (STL) specifications at runtime without retraining, while remaining agnostic of the underlying foundation model. We perform comprehensive evaluation of our approach across state-of-the-art navigation foundation models and we show that our decoding-time interventions are useful not only for filtering unsafe actions but also for conditional action-generation. Videos and code are available on our website: https://constrained-robot-fms.github.io
@@ 1 Introduction
Recent advances in Robotics Foundation Models (RFM) have enabled general purpose robot policies that map multi-modal inputs such as RGB images, natural language instructions and proprioceptive inputs to action sequences [1]. RFMs such as SPOC [2], PoliFormer [3], OpenVLA [4] exhibit impressive generalization in navigation and manipulation tasks and serve as versatile robot controllers for real-world deployment contexts. However, these models are primarily data-driven and lack any explicit notion of safety. Although these models may implicitly exhibit safety-related behaviors depending on the patterns in their training data, there is no formal guarantee that models will consistently behave safely in all situations. This serves as a limiting factor for deploying these foundation models in the physical world where rule compliance and regulatory safety rule adherence are crucial.
@@
Formal specifications have long been used to specify rules of operation such as safety requirements or mission directives for robotic deployments [5, 6]. Specifically, temporal logics [7]  can capture rich desired temporal constraints over robot behavior, such as “remain within permitted region zones and avoid dangerous obstacles”. This provides a rigorous workflow to encode behavior rulebooks using temporal logics and enforce them for robotic foundation models. Although temporal logic has seen success in classical robotic planning for reliable behavior generation, its use for foundation models remains limited. Additionally, retraining or fine-tuning these large pre-trained models to directly embed temporal logic specification is challenging [8]. First, retraining models is costly endeavor in terms of computational resources and data requirements. Moreover, due to the stochastic nature of these models, it is difficult to guarantee strict satisfaction of temporal constraints through training alone. Hence, there is a pressing need for methods that can enforce temporal specifications efficiently at inference time without disrupting the model’s pre-trained behavior.
@@
In the field of natural language processing, syntactic constraints have been successfully enforced by applying constrained decoding at inference time [9, 10, 11]. These approaches typically mask out tokens that violate a syntactic constraint defined over token sequences. For example, regular expressions (regex) represent a widely used form of syntactic constraint, requiring that generated token sequences conform to predefined structural patterns [9, 10]. Inspired by this line of work, we extend the paradigm of constrained decoding to enforce logical constraints over action trajectories in dynamical systems and propose specification aligned decoding (SpecDec) for RFMs that ensures generated action sequences provably satisfy Signal Temporal Logic (STL) [12] specifications. Our key insight is that decoding-time interventions can be used not just to filter unsafe actions, but to condition the generation process itself on specification satisfaction. This conditioning is critical because it steers the model toward generating specification satisfying actions rather than relying on post hoc rejection. SpecDec reduces risk of infeasible outputs while preserving the original action distribution of the model. To enforce such specifications, we leverage the formal semantics of STL to evaluate candidate actions at runtime and mask those that lead to future violations. Our method is agnostic to the underlying foundation model, requiring only two properties: (1) access to the decoding-layer logits during inference, and (2) access to a lightweight dynamics model to predict future states. To efficiently evaluate STL specifications at inference time, we use a high-performance computational graph based library STLCG++ [13]. To the best of our knowledge, this is the first work to apply constrained decoding with STL specifications to RFMs.
@@
Our main contributions are as follows: First, we introduce the novel problem of constrained decoding for robotic foundation models under Signal Temporal Logic specifications (Section 3.1). Second, we propose an inference-time technique that reweights or masks candidate actions using STL satisfaction scores in (Section 3.2 and 3.3). Finally, we demonstrate the effectiveness of our method on state-of-the-art object navigation models without modifying model parameters (Section 4).
@@ 2.1 Signal Temporal Logic
Apart from the Boolean semantics, quantitative semantics are defined for a signal to compute a real-valued metric indicating robustness, i.e., the strength of satisfaction or violation. For the sake of brevity, the definition of robustness is provided in Appendix A.
@@ 3 Specification-Guided Constrained Decoding
Figure 1: Overview of our specification aligned decoding framework. Given multimodal inputs (e.g., RGB images and natural language instructions), a pretrained Robot Foundation Model (e.g., SPOC) generates candidate actions. These actions are filtered or reweighted by the constrained decoding technique based on temporal logic constraints. In the figure, green markers denote target object locations, while red zones represent regions to avoid, as defined by temporal logic constraints.
In this section, we introduce a novel problem formulation for SpecDec in RFMs. First, we highlight the challenge in specification checking for RFMs in contrast to traditional syntactical constraint checking adopted by LLMs, and our solution to remedy it. Then, we propose two novel inference-time techniques for specification aligned decoding: Hard Constrained Decoding (HCD) and Robustness Constrained Decoding (RCD).
@@ 3.1 Problem Statement
As highlighted in the background section, existing techniques in constrained decoding for language models enforce syntactic constraints defined over tokens such as conforming to a context-free grammar or matching a regular expression. In these setups, constraint checking can be performed in the model’s token space.
In contrast, RFMs operate in a physical environment and constraints (captured via temporal logic) are defined over state variables. Since a large class of end-to-end RFMs solely propose action sequences, specification checking can only be performed as actions are executed and the environment is simulated forward. In this case, constrain checking cannot be done solely in the token space and requires environmental feedback or a dynamics stepping function. Hence, we leverage a dynamics function to compute specification satisfaction of different action sequences proposed by an RFM.
@@
In this work, we focus on RFMs that generate actions based on multi-modal inputs, including sensor observations (e.g., RGB, depth, LiDAR) and natural language instructions. Let $\mathcal{I}_{t}$ represent the aggregated input at timestep $t$. These inputs are first encoded into a latent embedding space through modality-specific encoders: $e_{\mathcal{I}_{t}}=\mathcal{E}_{\mathcal{I}}\!\bigl{(}\mathcal{I}_{t}\bigr{)}.$
Given the history of encoded inputs up to time $t$, a Transformer-based foundation model parameterized by $\theta$ predicts embeddings for the next TtT{-}t actions:
@@
Most of the existing techniques for specification enforcement perform posthoc manipulation of proposed foundation model actions through filtering or rejecting action sequences that violate the specification $\varphi$. Although manipulation after sampling can ensure specification satisfaction, it can lead to distorting the model’s learned distribution, producing low likelihood outputs. This undermines the inductive biases learned during pretraining and leads to degenerate, brittle behaviors. A similar problem was highlighted when ensuring compliance with logical constraints for large language models in [20]. Additionally, RFMs decode actions sequentially, where each action $a_{t}$ is conditioned on previously generated tokens $a_{<t}$. Posthoc manipulation can disrupt this causal chain and lead to a mismatch between the model’s internal hidden state and the executed sequence. Hence, we propose the following problem statement:
@@
Let $\pi(a_{1:T})$ be the unconstrained action-sequence distribution produced by the RFM’s decoder (e.g. the softmax over logits generated by the Transformer). We define the ideal constrained distribution over action sequences as:
$Q_{\pi,\varphi}(a_{1:T})=\frac{\pi(a_{1:T})\cdot\mathbf{1}[(x_{1:T},a_{1:T})\models\varphi]}{\sum\limits_{a_{1:T}^{\prime}}\pi(a_{1:T}^{\prime})\cdot\mathbf{1}[(x_{1:T}^{\prime},a_{1:T}^{\prime})\models\varphi]}$ (1)
where $x_{1:T}$ denotes the state trajectory induced by the system dynamics under actions $a_{1:T}$ and $\mathbf{1}[\cdot]$ is the indicator function that returns 1 iff the trajectory-action pair satisfies the specification. Equation 1 is the exact Bayesian conditioning of $\pi$ on the event that the generated rollout satisfies $\varphi$. Hence, sampling from $Q_{\pi,\varphi}$ would give sequences that (i) inherit the original model’s preferences encoded in $\pi$ and (ii) guarantee specification satisfaction.
In this work, we propose a technique to overcome the drawbacks of post-hoc safety enforcement methods (such as filtering) by leveraging constrained decoding techniques. Specifically, we propose SpecDec: A constrained decoding strategy that integrates STL specifications into the foundation model action selection process itself, ensuring satisfaction without distorting the model’s distribution.
@@ 3.2 Hard Constrained Decoding
Here $t$ is the current decision step, $k$ is an index for the look-ahead step $t+k$ within a planning horizon of length $T$ ($k\in[1..T]$), $\hat{a}_{t+k}^{(i)}$ is the action mapping to the token i and $\hat{x}_{t+k}^{(i)}$ is the next state value upon taking this action. This next state is elicited using a simple dynamics model ($f$) as highlighted in the previous section. Adjusting logits in this fashion ensures that any invalid token with respect to the safety specification will have zero probability of being selected after applying the softmax function. We introduce a theorem to state that our proposed approach generates a trajectory that satisfies the given specification $\varphi$:
Theorem 1 (Specification Satisfaction Modulo Dynamics).
Let $\pi_{\text{HCD}}$ denote the policy induced by hard-constrained decoding with respect to an STL specification $\varphi$, and let $f$ be a deterministic dynamics model used during decoding. Then the generated trajectory satisfies the specification:
$\{(x_{t+1},\hat{a}_{t+1}),\dots,(x_{T},\hat{a}_{T})\}\models\varphi$
where each action $\hat{a}_{t+k}\sim\pi_{\text{HCD}}(\cdot\mid x_{t+k-1})$, and $x_{t+k}=f(x_{t+k-1},\hat{a}_{t+k})\quad\text{for all }k.$
A proof sketch is provided in the Appendix B
@@ 3.3 Robustness Constrained Decoding
HCD ensures compliance but can lead to compromising task success, which can be undesirable. A similar tradeoff was observed by [21] when probability space-steering preserved model fluency while reducing toxic continuations compared with hard-filtering strategies that inflated perplexity and eroded diversity. Hence, we propose an alternative approach, called RCD, where we leverage the quantitative semantics of STL specifications (robustness). Unlike HCD, which applies hard masking to completely remove unsafe actions, RCD softly guides the model toward safer actions by incorporating robustness scores that reflect the degree of satisfaction of $\varphi$. This is similar to the approach proposed in [21] where the next-token distributions were re-weighted based on the utility scores provided by another language model. Our utility scores are quantified by the robustness function( $\rho(x_{t},\varphi)$ ) that returns a real-valued score indicating how well a predicted state satisfies the specification. Positive robustness values denote specification satisfaction, while negative values capture the degree of violation.
@@
This approach allows for graded preferences that improve flexibility and robustness to dynamics approximation errors. This is because every action keeps a non-zero probability, the policy can still recover when the dynamics model is imperfect. Concretely, if the predicted successor $\hat{x}_{t+1}$ is off by $\epsilon$, an action that looked marginally unsafe can be safe in the true system, and vice versa. Retaining a weighted down probability for this action gives the sampler a fall-back option where as HCD would completely rule this action out due to 0 probability. Since we are shifting the probability mass for unsatisfying actions, it is possible that they are still chosen and lead to a violation. However, this is a tradeoff we allow to achieve a given task objective. We note that this still ensures higher STL sastifaction than unconstrained actions.
@@ 4.1 Implementational Details
We evaluate our constrained decoding approach on AI2-THOR simulated [22] indoor environments using the state-of-the-art (SOTA) navigation model called Shortest Path Oracle Clone (SPOC) [2]. This model achieves high navigation success rates for diverse tasks such as object navigation, room navigation, and so forth. We encode the STL specifications using an efficient computational graph-based STL library called STLCG++ that can evaluate multiple state signals in parallel [13]. This ensures minimal inference overhead at runtime, which is crucial for foundation model deployment. We enforce geofencing and obstacle avoidance by encoding them into invariant STL safety specifications. Specifically, we generate random regions in the configuration space that the robot must either avoid (obstacle zones) or remain within (safe zones), and apply these constraints in real time during execution.
4.2 Benchmarks and Research Questions

### v3

@@ Abstract
Recent advances in the development of robotic foundation models have led to promising end-to-end and general-purpose capabilities in robotic systems. Trained on vast datasets of simulated and real-world trajectories, these models map multimodal observations directly to action sequences for physical execution. Despite promising real-world capabilities, these models are still data-driven and, therefore, lack explicit notions of behavioral correctness. We address this gap by introducing SafeDec, a constrained decoding framework for autoregressive, robot foundation models that enforces invariant safety specifications on candidate action trajectories. Task-specific safety rules are expressed as Signal Temporal Logic (STL) formulas and are enforced at inference time with minimal overhead. Our method ensures that generated actions provably satisfy STL specifications under assumed dynamics at runtime without retraining , while remaining agnostic of the underlying policy. We evaluate SafeDec on tasks from the CHORES benchmark for state-of-the-art generalist policies (e.g., SPOC, Flare, PoliFormer) across hundreds of procedurally generated environments and show that our decoding-time interventions are useful not only for filtering unsafe actions but also for conditional action generation. Videos are available at constrained-robot-fms.github.io.
@@ 1 Introduction
Recent advances in developing Robotics Foundation Models (RFM) have enabled general-purpose policies that map multi-modal inputs such as RGB images, natural language instructions, and proprioceptive inputs to action sequences Hu et al. [1]. Shortest Path Oracle Clone (SPOC) [2], PoliFormer [3], Flare [4] and OpenVLA [5] exhibit impressive generalization in navigation and manipulation tasks and serve as versatile robot controllers for real-world deployment contexts. However, these models are primarily data-driven and lack any explicit notion of safety. Although these models may implicitly exhibit safety-related behaviors depending on the patterns in their training data, there is no formal guarantee that models will consistently behave safely in all situations. This serves as a limiting factor for deploying these foundation models in the physical world where rule compliance and regulatory safety rule adherence are crucial.
@@
Formal specifications have long been used to specify safety requirements for robotic deployments [6, 7]. Temporal logics (TL) [8]  can capture safety constraints on robot behavior, such as “remain within the permitted region zones and avoid dangerous obstacles”. Although TL has seen success in classical robotic planning for safety constraint satisfaction, its use for enforcing safety for large transformer-based robot policies remains limited. Additionally, retraining or fine-tuning these large pre-trained models to directly embed temporal logic specification is challenging [9]. First, retraining models is a costly endeavor in terms of computational resources and data requirements. Moreover, due to the stochastic nature of these models, it is difficult to guarantee strict satisfaction of safety constraints through training alone. Hence, there is a pressing need for methods that can enforce safety specifications efficiently at inference time without disrupting the model’s pre-trained behavior.
@@
In the field of natural language processing, syntactic constraints have been successfully enforced by applying constrained decoding at inference time [10, 11, 12]. These approaches typically mask out tokens that violate a syntactic constraint defined over token sequences. For example, regular expressions (regex) represent a widely used form of syntactic constraint, requiring that generated token sequences conform to predefined structural patterns [10, 11]. Inspired by this line of work, we extend the paradigm of constrained decoding to enforce safety constraints over action trajectories in dynamical systems and propose safety specification aligned decoding (SafeDec) for transformer based policies that ensures generated action sequences provably satisfy Signal Temporal Logic (STL) [13] specifications under assumed dynamics. Our key insight is that decoding-time interventions can be used not just to filter unsafe actions, but to condition the generation process itself on specification satisfaction. This conditioning is critical because it steers the model toward generating safety specification satisfying actions rather than relying on post hoc rejection. SafeDec reduces risk of infeasible outputs while preserving the original action distribution of the model. To enforce such specifications, we leverage the formal semantics of STL to evaluate candidate actions at runtime and mask those that lead to future violations. Our method is agnostic to the underlying foundation model, requiring only two properties: (1) access to the decoding-layer logits during inference, and (2) access to an approximate dynamics model to predict future states. To efficiently evaluate STL specifications at inference time, we use a high-performance computational graph based library STLCG++ [14]. To the best of our knowledge, this is the first work to enforce formal safety on transformer-based robotic policies at inference time using constrained decoding.
@@
Our main contributions are as follows: First, we introduce the novel problem of constrained decoding for transformer-based policies under safety STL specifications (Section 3.1). Second, we propose an inference-time technique that reweights or masks candidate actions using STL satisfaction scores in (Section 3.2 and 3.3). Finally, we demonstrate the effectiveness of our method on state-of-the-art object navigation models without modifying model parameters (Section 4).
@@ 2.1 Signal Temporal Logic
Apart from the Boolean semantics, quantitative semantics are defined for a signal to compute a real-valued metric indicating robustness, i.e., the strength of satisfaction or violation. For the sake of brevity, the definition of robustness is provided in Appendix A.1.
@@ 3 Specification-Guided Constrained Decoding
Figure 1: Overview of our specification aligned decoding framework. Given multimodal inputs (e.g., RGB images and natural language instructions), a pretrained robot foundation model (e.g., SPOC) generates candidate actions. These actions are filtered or reweighted by the constrained decoding technique based on STL safety constraints. In the figure, green markers denote target object locations, while red zones represent regions to avoid.
In this section, we introduce a novel problem formulation for SafeDec in RFMs. First, we highlight the challenge in specification checking for these models in contrast to traditional syntactical constraint checking adopted by LLMs, and our solution to remedy it. Then, we propose two novel inference-time techniques for specification aligned decoding: Hard Constrained Decoding (HCD) and Robustness Constrained Decoding (RCD).
@@ 3.1 Problem Statement
As highlighted in the background section, existing techniques in constrained decoding for language models enforce syntactic constraints defined over tokens, such as conforming to a context-free grammar or matching a regular expression. In these setups, constraint checking can be performed in the model’s token space.
In contrast, RFMs operate in a physical environment and constraints (captured via temporal logic) are defined over state variables. Since a large class of end-to-end policies solely propose action sequences, specification checking can only be performed as actions are executed and the environment is simulated forward. In this case, constraint checking cannot be done solely in the token space and requires environmental feedback or a dynamics stepping function. Hence, we leverage an approximate first order dynamics function to compute specification satisfaction of different action sequences proposed by these policies.
@@
In this work, we focus on RFMs that generate actions based on multi-modal inputs, including sensor observations (e.g., RGB, depth, LiDAR) and natural language instructions. Let $\mathcal{I}_{t}$ represent the aggregated input at timestep $t$. Given the history of encoded inputs up to time $t$, a transformer-based policy parameterized by $\theta$ predicts embeddings for the next TtT{-}t actions:
@@
Most of the existing techniques for specification enforcement perform posthoc manipulation of proposed actions through filtering or rejecting action sequences that violate the specification $\varphi$. Although manipulation after sampling can ensure specification satisfaction, it can lead to distorting the model’s learned distribution, producing low-likelihood output. This undermines the inductive biases learned during pretraining and leads to degenerate, brittle behaviors. A similar problem was highlighted when ensuring compliance with logical constraints for large language models (LLM) in Park et al. [21]. Additionally, these models decode actions sequentially, where each action $a_{t}$ is conditioned on previously generated tokens $a_{<t}$. Posthoc manipulation can disrupt this causal chain and lead to a mismatch between the model’s internal hidden state and the executed sequence. Hence, we propose the following problem statement:
@@
Let $\pi(a_{0:T})$ be the unconstrained action-sequence distribution produced by the transformer-based policy’s decoder (e.g. the softmax over logits generated by the Transformer). We define the ideal constrained distribution over action sequences as:
$Q_{\pi,\varphi}(a_{0:T})=\frac{\pi(a_{0:T})\cdot\mathbf{1}[(x_{1:T},a_{0:T})\models\varphi]}{\sum\limits_{a_{0:T}^{\prime}}\pi(a_{0:T}^{\prime})\cdot\mathbf{1}[(x_{1:T}^{\prime},a_{0:T}^{\prime})\models\varphi]}$ (1)
where $x_{1:T}$ denotes the state trajectory induced by the system dynamics under actions $a_{0:T}$ and $\mathbf{1}[\cdot]$ is the indicator function that returns 1 iff the trajectory-action pair satisfies the specification. Equation 1 is the exact Bayesian conditioning of $\pi$ on the event that the generated rollout satisfies $\varphi$. Hence, sampling from $Q_{\pi,\varphi}$ would give sequences that (i) inherit the original model’s preferences encoded in $\pi$ and (ii) guarantee specification satisfaction.
In this work, we propose a technique to overcome the drawbacks of post-hoc safety enforcement methods (such as filtering) by leveraging constrained decoding techniques. Specifically, we propose SafeDec : A constrained decoding strategy that integrates STL specifications into the foundation model action selection process itself, ensuring satisfaction without distorting the model’s distribution.
@@ 3.2 Hard Constrained Decoding
Here $t$ is the current decision step, $k$ is an index for the look-ahead step $t+k$ within a planning horizon of length $T$ ($k\in[1..T]$), $\hat{a}_{t+k}^{(i)}$ is the action mapping to the token i and $\hat{x}_{t+k}^{(i)}$ is the next state value upon taking this action. This next state is elicited using a simple dynamics model ($f$) as highlighted in the previous section. Adjusting logits in this fashion ensures that any invalid token with respect to the safety specification will have zero probability of being selected after applying the softmax function.
@@ 3.3 Robustness Constrained Decoding
HCD ensures compliance but can lead to compromising task success, which can be undesirable. A similar tradeoff was observed by Liu et al. [22] when probability space-steering preserved model fluency while reducing toxic continuations compared with hard-filtering strategies that inflated perplexity and eroded diversity. Hence, we propose an alternative approach, called RCD, where we leverage the quantitative semantics of STL specifications (robustness). Unlike HCD, which applies hard masking to completely remove unsafe actions, RCD softly guides the model toward safer actions by incorporating robustness scores that reflect the degree of satisfaction of $\varphi$. This is similar to the approach proposed in Liu et al. [22] where the next-token distributions were re-weighted based on the utility scores provided by another language model. Our utility scores are quantified by the robustness function( $\rho(x_{t},\varphi)$ ) that returns a real-valued score indicating how well a predicted state satisfies the specification. Positive robustness values denote specification satisfaction, while negative values capture the degree of violation. A formal definition of robustness in line with the STL quantitative semantics is provided in Appendix A.1.
@@
This approach allows for graded preferences that improve flexibility and robustness to dynamics approximation errors. By ensuring that all actions preserve a non-zero probability, the policy remains capable of recovering from errors arising from an imperfect dynamics model. Concretely, if the predicted successor $\hat{x}_{t+1}$ is off by $\epsilon$, an action that appeared marginally unsafe can be safe in the true system, and vice versa. Retaining a weighted down probability for this action gives the sampler a fall-back option whereas HCD would completely rule this action out due to 0 probability. Since we are shifting the probability mass for unsatisfying actions, it is possible that they are still chosen and lead to a violation. However, this is a tradeoff that we allow to achieve a given task objective. We note that this still ensures higher STL satisfaction than unconstrained actions.
@@ 4.1 Implementational Details
We comprehensively evaluate our constrained decoding framework on procedurally generated AI2-THOR [23] indoor scenes with diverse objects and layouts using three state-of-the-art (SOTA) RFMs: Shortest Path Oracle Clone (SPOC) [2], PoliFormer [3] and Flare [4]. All three are large transformer-based embodied agents trained on extensive language-conditioned robot trajectory datasets. These models achieve strong zero-shot generalization for a vast variety of navigation tasks that span open vocabulary object-goal navigation (“find a mug”), room-to-room traversal (“visit all rooms”), waypoint-based navigation (“move three meters forward and stop near the red rug”), and attribute-conditioned variants (“locate the chair closest to the refrigerator in the kitchen”). These models also demonstrate reliable zero-shot transfer to real-world environments, achieving robust task satisfaction.
In addition, these models capture three different training paradigms for RFMs. SPOC is trained purely with imitation learning from shortest-path rollouts. Poliformer employs a hybrid approach that combines reinforcement learning and imitation learning, enabling it to learn long-horizon structure while retaining expert priors. Flare adopts a large-scale pretraining plus fine-tuning on embodied navigation data in line with recent foundation model training paradigms. This diversity in training paradigms allows us to evaluate the applicability of SafeDec across different learning regimes.
In this work, we address safety specifications for robotics and, therefore, select those most relevant to real-world deployment. In particular, we focus on an important class of safety specifications called invariants, which are specifications that must be enforced at every reachable state of the system (e.g., “always avoid an unsafe region”). We enforce geofencing and obstacle avoidance by encoding them as invariant specifications in STL. Specifically, we generate random regions in the configuration space that the robot must either avoid (obstacle zones) or remain within (safe zones), and apply these constraints in real time during execution. The specifications used are: $\varphi_{\text{geofence}}=\mathbf{G}\!\left(\bigvee_{i=1}^{N}\Bigl(x_{i}^{\mathrm{L}}\leq x\leq x_{i}^{\mathrm{U}}\;\wedge\;z_{i}^{\mathrm{L}}\leq z\leq z_{i}^{\mathrm{U}}\Bigr)\right)$, $\varphi_{\text{avoid}}\;=\;\mathbf{G}\!\left(\bigwedge_{i=1}^{N}\neg\Bigl(x_{i}^{\mathrm{L}}\leq x\leq x_{i}^{\mathrm{U}}\;\wedge\;z_{i}^{\mathrm{L}}\leq z\leq z_{i}^{\mathrm{U}}\Bigr)\right)$. The size of the regions for $\varphi_{avoid}$ is 1 $m^{2}$. For $\varphi_{geofence}$, we randomly pick a subset of rooms in each house and use each chosen room’s full bounds. These safety specifications are common for robot learning applications [24, 25, 26, 27]. We encode our test STL specifications using an efficient computational graph-based STL library called STLCG++ that can evaluate multiple state signals in parallel [14]. This ensures minimal inference overhead at runtime ($10^{-5}$ s per timestep) , which is crucial for policy deployment. For our dynamics model, we assume a unicycle model, an approximate first-order dynamics abstraction widely used in the robotics literature for analysis and control [28]. This representation captures the essential kinematics of motion in the plane and is widely used because it is applicable for diverse robotic platforms.
4.2 Experimental setup